OpenStudy (andertj21):

The first side of a triangle measures 4 in. less than the second side, the third side is 3 in. more than the first side, and the perimeter is 15 in. How long is the third side? If s represents the length of the second side, which of the following represents the length of the third side? s - 4 s - 1 s + 3

OpenStudy (aric200):

s = length of second side s - 4 = length of first side s - 4 + 3 = s - 1 = length of the third side s - 1

OpenStudy (andertj21):

Thanks dude